# HOT3D: Hand and Object Tracking in 3D from Egocentric Multi-View Videos

### v1

@@ 4 Experiments
Figure 9: Example results of 3D lifting of in-hand objects. Ground-truth 3D object locations are shown in green, locations predicted by StereoMatch in red, by MonoDepth in blue, and by HandProxy in orange. The 3D locations are projected to the three views from Aria. In each sample, the RGB image is on the left and monochrome images on the right. The DINOv2 stereo matching between the RGB and one of the monochrome Aria images is visualized in the top left corners.
Results (Table 5) The StereoMatch method outperforms the MonoDepth method on Aria recordings, regardless of whether the ground-truth segmentation masks or masks predicted by MRCNN-DA are used as input. The 3D localization error of MonoDepth is primarily oriented along the optical axis. This is evident in Fig. 9, where the localization error for the MonoDepth approach is most pronounced in camera viewpoints with a larger baseline w.r.t. the input RGB camera. Finally, the HandProxy baseline demonstrates that while hand tracking alone is sufficient for localizing objects within a 30 cm uncertainty radius (similar to evaluations in [51]), a dedicated method for multi-view 3D lifting of in-hand objects provides additional value for finer-grained localization within 10 cm.
@@ 5 Conclusion
We have introduced HOT3D, a large-scale, publicly available dataset designed to support training and evaluation of methods for various 2D and 3D egocentric tasks related to hand-object interaction. Our experiments show that multi-view methods, whose benchmarking is uniquely enabled by HOT3D, significantly outperform their single-view counterparts across several popular tasks. Besides multi-view 3D object tracking, we addressed the tasks of multi-view 6DoF pose estimation of unseen objects and 3D lifting of in-hand objects, for which we developed strong baselines. By publicly releasing the HOT3D dataset along with source code of the baseline methods, and by co-organizing public challenges [^1] https://bop.felk.cvut.cz/challenges/bop-challenge-2024/ [^2] https://github.com/facebookresearch/hand_tracking_toolkit on the dataset, we hope to accelerate research on egocentric vision and contextual AI.
Acknowledgements. Thanks to Ben Schneiders, Mateja Kodila, Elizabeth Béres, and Nicolas Greiner for organizing data collections and helping with data processing. Thanks to Fadime Sener, Bugra Tekin, Yann Labbé, Christian Forster, Mariano Jaimez, Rainer Stal, Timo Stoffregen, Pierluigi Taddei, Eric Sauser, Shitai Li, Lukas Bode, and Ben Schneiders for participating in the data collections. Thanks to Evgeniy Oleinik and Maien Hamed for help with hosting the dataset on Meta servers. Thanks to Xiaqing Pan for his guidance on open sourcing the HOT3D toolkit. Thanks to Mark Schwesinger, Christopher Pistritto, Adam Szenderski, Elizabeth Argall, and Josiah Zacharias for providing legal and writing support to release HOT3D on time. Thanks to Nick Charron, Alexander Gamino, and David Caruso for their help with calibration and SLAM-OptiTrack alignment. Thanks to Ka Chen and Allison Tilp for supporting the creation of high-fidelity CAD models of the HOT3D objects. Thanks to Abha Arora, Luis Pesqueira, and Yuyang Zou for helping with organizing the Aria data campaign and setting up the Aria data collection infrastructure. Thanks to Tomasz Malisiewicz for his help with training models for 2D in-hand object segmentation.
@@ Appendix A Aria glasses
Project Aria [13] is an egocentric recording device in glasses form-factor created by Meta. It is designed as a research tool for egocentric machine perception and contextualized AI research, and available to researchers across the world via projectaria.com.
@@ A.1 Device and sensors
Project Aria is built to emulate future AR- or smart-glasses catering to machine perception and egocentric AI rather than human consumption. It is designed to be wearable for long periods of time without obstructing or impeding the wearer, allowing for natural motion even when performing highly dynamic activities – such as playing soccer or dancing. It has a total weight of 75g (compared to over 150g for a single GoPro camera), and fits just like a pair of glasses.
@@
  • One rolling-shutter RGB camera recording at 30 fps and $1408\times 1408$ resolution. It is fitted with an F-Theta fisheye lens that covers a field of view of $110^{\circ}$.
  • Two global-shutter monochrome cameras recording at 30 fps and $640\times 480$ resolution. They provide additional peripheral vision, and are fitted with F-Theta fisheye lenses that cover a field of view of $150^{\circ}$.
  • Two monochrome eye-tracking cameras recording at 10 fps and $320\times 240$ resolution.
@@ A.3 Processing summary
Then, the MPS pipeline is invoked for each full Aria recording, which are typically about 2 minutes long and include many instances of hand-object interactions with different objects. Next, we 7DoF-align the MPS output with the OptiTrack coordinate frame (App. C). In total, we have processed 199 Aria recordings with a total length of 391 minutes.
@@ A.4 Tools and ecosystem
Object in hand (mIoU $\uparrow$) for training + test / training / test split:
@@
Recall [%] $\uparrow$ for training + test / training / test split:
@@
Appendix B Quest 3 headset
Quest 3 [41], shown in Fig. 13, is the latest production headset from Meta for virtual- and mixed-reality experiences. For the HOT3D data collection we used a developer version of the Quest 3 headset. This version has four global-shutter monochrome cameras with fisheye lenses, 1280x1024 px image resolution, 18 PPD (Pixels Per Degree), and records at 30 fps. Two of the cameras are on the front side of the headset, roughly aligned with eyes, and two on the sides. HOT3D only includes images from the two front cameras as those capture the relevant scene part (the two side cameras are useful for applications like SLAM). Example images are in Fig. 14. Data from other sensors present in the consumer version of Quest 3, including a gyroscope and an accelerometer, were not recorded. The intrinsic and extrinsic parameters of the headset cameras were calibrated with a ChArUco board. Both the headset and the board were attached a set of optical markers and tracked by the motion-capture system described in App. C, which allowed to estimate camera-to-headset transformations. At recording time, the headset pose was still tracked by the motion-capture system and used to calculate per-frame camera-to-world transformations.
[…]
Appendix C Marker-based motion capture
Appendix D Object orientation statistics
When recording HOT3D, we asked subjects to naturally interact with the objects. Consequently, orientation distributions of the observed objects (Fig. 15) reveal clear object-specific pose biases, which may be useful as prior information at inference time (we see that the bowl tends to be seen upright, the birdhouse from the front and upright, etc.).
Appendix E Additional quantitative results
The results of 2D segmentation and 3D lifting of in-hand objects presented in Tables 4 and 5 were obtained by evaluating methods on clips from both training and test splits. To allow the community to compare their results against our results on these two tasks, we additionally provide results obtained on clips from the training split for which the ground-truth annotations are publicly available (Tables 6 and 7). Evaluating on the training split is possible as both of these tasks are training-free and therefore do not require any training split.

### v2

@@ 4 Experiments
Figure 9: Example results of 3D lifting of in-hand objects. Ground-truth 3D object locations are shown in green, predictions from StereoMatch in red, MonoDepth in blue, and HandProxy in orange. The locations are projected to the three Aria views. In each example, the RGB image is on the left, monochrome images on the right, and DINOv2 stereo matching between the RGB and a monochrome image in the top left.
4.4 3D lifting of in-hand objects
Experimental setup. Finally, we focus on the task of 3D lifting of unknown in-hand objects, which is useful for object indexing and long-term tracking [51]. Given per-view 2D segmentation masks of an in-hand object, the goal is to estimate the 3D object location. We developed and compare the performance of three methods described below. The accuracy of the estimated locations is measured by a recall rate defined as the fraction of samples for which a correct location was estimated. A location is considered correct if its offset from the ground-truth location, defined by the center of the 3D object bounding box, is below a threshold.
Using hand as the object proxy (HandProxy). As a simple baseline, we use the ground-truth 3D palm center as the 3D object location. We define the palm center as the middle point between the wrist joint and the first joint of the middle finger. If both hands are visible, we use the 3D palm location whose 2D projection is closer to the 2D object mask centroid. The goal of this baseline is to evaluate whether a specialized solution is necessary for this task or whether an accurate 3D hand tracker could provide a sufficient estimate.
Lifting by monocular depth estimation (MonoDepth). Our second approach is inspired by [51] and relies on monocular depth prediction and sparse SLAM observations. Specifically, we predict a monocular depth map by applying Depth Anything V2 [72] to a rectilinear input image (warped from the fisheye camera to a pinhole camera), align it to the 3D point cloud from SLAM by a scale-shift transformation [51], and finally calculate the 3D object location as the median of 3D points given by the aligned depth values collected from the object mask. Since MonoDepth requires SLAM observations that are not available for Quest 3, we evaluate this method only on recordings from Aria (using RGB images).
Results (Tab. 5). The StereoMatch method outperforms MonoDepth regardless of whether the ground-truth segmentation masks or masks predicted by MRCNN-DA are used as input. The 3D localization error of MonoDepth is primarily along the optical axis and therefore caused by inaccurate depth predictions. This is evident in Fig. 9 where the localization error is most pronounced in camera viewpoints with a larger baseline w.r.t. the input RGB camera. Finally, the HandProxy baseline demonstrates that while hand tracking alone is sufficient for localizing objects within a 30 cm uncertainty radius (similar to evaluations in [51]), a dedicated method for multi-view 3D lifting of in-hand objects provides additional value for finer-grained localization within 10 cm.
@@ 5 Conclusion
We have introduced HOT3D, a large-scale dataset designed to facilitate the training and evaluation of methods for various 2D and 3D egocentric tasks related to hand-object interaction. Our experiments show that multi-view methods, whose benchmarking is uniquely enabled by HOT3D, significantly outperform their single-view counterparts across several popular tasks. Besides multi-view 3D object tracking, we addressed the tasks of multi-view 6DoF object pose estimation and 3D lifting of in-hand objects, for which we developed strong baselines. By publicly releasing the dataset and co-organizing public challenges [^1] https://bop.felk.cvut.cz/challenges/bop-challenge-2024/ [^2] https://github.com/facebookresearch/hand_tracking_toolkit on the dataset, we hope to accelerate research on egocentric vision and contextual AI.
Acknowledgements
Big thanks to Lingni Ma for the initial discussions, insightful ideas, and sharing her expertise with motion capture, which all served as the foundation for the HOT3D project. Thanks to Kevin Harris and Steve Olsen for their invaluable expertise and help with setting up the motion-capture lab. Thanks to Ben Schneiders, Mateja Kodila, Elizabeth Béres, Pablo Gomez Ponce, and Nicolas Greiner for organizing data collections and helping with data processing. Thanks to Fadime Sener, Bugra Tekin, Yann Labbé, Christian Forster, Mariano Jaimez, Rainer Stal, Timo Stoffregen, Pierluigi Taddei, Eric Sauser, Shitai Li, Lukas Bode, and Ben Schneiders for participating in the data collections. Thanks to Evgeniy Oleinik and Maien Hamed for help with hosting the dataset on Meta infrastructure. Thanks to Xiaqing Pan for his guidance on open sourcing the HOT3D toolkit. Thanks to Mark Schwesinger, Christopher Pistritto, Elizabeth Argall, and Josiah Zacharias for providing legal and writing support to release HOT3D on time. Thanks to Nick Charron, Alexander Gamino, and David Caruso for their help with calibration and SLAM-OptiTrack alignment. Thanks to Ka Chen, Allison Tilp, and Thibaud Gayraud for supporting the creation of high-fidelity CAD models of the HOT3D objects. Thanks to Abha Arora, Luis Pesqueira, and Yuyang Zou for helping with organizing the Aria data campaign and setting up the Aria data collection infrastructure. Thanks to Tomasz Malisiewicz for his help with training models for 2D in-hand object segmentation. Thanks to Weiguang Si for his help with preparing HOT3D for the Hand Tracking Challenge at ECCV 2024.
Appendix
In this appendix, we provide details about the Aria glasses (Sec. A) and the Quest 3 headset (Sec. B) which were used to record the HOT3D dataset. We also describe our procedure for ground-truth annotation (Sec. C), provide additional data statistics (Sec. D) and quantitative results (Sec. E).
@@ Appendix A Aria glasses
Project Aria [13] is an egocentric recording device in glasses form-factor by Meta, designed as a research tool for egocentric machine perception and contextualized AI research, and available to researchers across the world via projectaria.com (Fig. 11).
@@ A.1 Device and sensors
Project Aria is built to emulate future AR/smart glasses catering to machine perception and egocentric AI rather than human consumption. Aria is designed to be wearable for long periods of time without obstructing or impeding the wearer, allowing for natural motion even when performing highly dynamic activities, such as playing soccer or dancing. Its total weight is 75g (a single GoPro camera has over 150g) and fits just like a pair of glasses.
@@
  • One rolling-shutter RGB camera recording at 30 fps and $1408\times 1408$ px. The camera is fitted with an F-Theta fisheye lens with a field of view (FOV) of $110^{\circ}$.
  • Two global-shutter monochrome cameras recording at 30 fps and $640\times 480$ px. These cameras provide additional peripheral vision and are fitted with F-Theta fisheye lenses with $150^{\circ}$ FOV.
  • Two monochrome eye-tracking cameras recording at 10 fps and $320\times 240$ px resolution.
@@ A.3 Processing summary
Then, the MPS pipeline is invoked for each full Aria recording, which are typically about 2 minutes long and include many instances of hand-object interactions with different objects. Next, we 7DoF-align the MPS output with the OptiTrack coordinate frame (Sec. C). In total, we have processed 199 Aria recordings with a total length of 391 minutes.
@@ A.4 Tools and ecosystem
Appendix B Quest 3 headset
Quest 3 [41], shown in Fig. 13, is the latest production headset from Meta for virtual- and mixed-reality experiences. For the HOT3D data collection we used a developer version of the Quest 3 headset. This version has four global-shutter monochrome cameras with fisheye lenses, 1280x1024 px image resolution, 18 PPD (Pixels Per Degree), and records at 30 fps. Two of the cameras are on the front side of the headset, roughly aligned with eyes, and two on the sides. HOT3D only includes images from the two front cameras as those capture the relevant scene part (the two side cameras are useful for applications like SLAM). Example images are in Fig. 14. Data from other sensors present in the consumer version of Quest 3, including a gyroscope and an accelerometer, were not recorded. The intrinsic and extrinsic parameters of the headset cameras were calibrated with a ChArUco board. Both the headset and the board were attached a set of optical markers and tracked by the motion-capture system described in Sec. C, which allowed to estimate camera-to-headset transformations. At recording time, the headset pose was still tracked by the motion-capture system and used to calculate per-frame camera-to-world transformations.
Appendix C Marker-based motion capture
@@
[…]
Appendix D Object orientation statistics
When recording, we asked subjects to naturally interact with the objects. Consequently, orientation distributions of the observed objects (Fig. 15) reveal clear object-specific pose biases, which may be useful as prior information at inference (we see that the bowl tends to be seen upright, the birdhouse from the front and upright, etc.).
Appendix E Additional quantitative results
The results of 2D segmentation and 3D lifting of in-hand objects presented in Tab. 4 and 5 were obtained by evaluating methods on clips from both training and test splits. To allow the community to compare their results against our results on these two tasks, in Tab. 6 and 7 we additionally provide results obtained on clips from the training split for which the ground-truth annotations are publicly available. Evaluating on the training split is possible as both of these tasks are training-free and therefore do not require any training split.
Object in hand (mIoU $\uparrow$) for training + test / training / test split
@@ Appendix E Additional quantitative results
Recall [%] $\uparrow$ for training + test / training / test split